OpenStudy (anonymous):

Determine the point(s) (if any) at which the graph of the function has a horizontal tangent line. (If an answer does not exist, enter DNE.) y = −6x + e^x

OpenStudy (mathmale):

Tangent line? How is that related to "derivative?"

OpenStudy (mathmale):

Does any of your learning material discuss how the derivative of a function represents the slope of the tangent line to the graph?

OpenStudy (mathmale):

Find the derivative with respect to x of y = −6x + e^x . The result, dy/dx, represents the slope of the tangent line to y = −6x + e^x. Please get started by finding this dy/dx.

OpenStudy (anonymous):

dy/dx=-6+e^x

OpenStudy (mathmale):

Yes. Now, at what point on the graph of y = −6x + e^x do you want to find the slope of the tangent line?

OpenStudy (mathmale):

Hint: Go back and re-read the original question.

OpenStudy (anonymous):

it doesn't give me any points But i have to find the zeros of this so i set the dy/dx = 0

OpenStudy (anonymous):

so -6+e^x = 0

OpenStudy (anonymous):

x = ln(6)?

OpenStudy (mathmale):

Pls explain further: why do you have to set the deriv. equal to zero?

OpenStudy (anonymous):

so that I can find the x coordinate and then I can plug that back into the original eqution

OpenStudy (mathmale):

"Determine the point(s) (if any) at which the graph of the function has a horizontal tangent line." First we want to know whether the graph actually has a horiz. tan. line (slope = 0). It does, because ... ????

OpenStudy (anonymous):

the correct anwser is (x,y) = ( ln(6), 6-6ln(6) ) which I have no idea how they got to that except for the ln (6)

OpenStudy (mathmale):

I was hoping you'd answer my question. How do we know that the graph of this function has a horiz. tangent line? In other words, do we know that the slope of the T. L. to this graph is zero anywhere?

OpenStudy (anonymous):

no I do not know if there is a zero anywhere

OpenStudy (mathmale):

Hint: The slope of the t. l. is given by dy/dx = -6 + e^x.

OpenStudy (mathmale):

In other words, do we know that the slope of the T. L. to this graph is zero anywhere?

OpenStudy (mathmale):

Answer: Yes. Question for you: Why?

OpenStudy (anonymous):

I do not know

OpenStudy (anonymous):

minimum?

OpenStudy (mathmale):

Because dy/dx = -6 + e^x =0 has a ..... ? (no, it's not minimum). Try just once more, please.

OpenStudy (mathmale):

From before: "In other words, do we know that the slope of the T. L. to this graph is zero anywhere?"

OpenStudy (mathmale):

The slope is dy/dx = -6 + e^x, and we have to set this = to 0 because of our search for a possible horiz. tan. line.

OpenStudy (mathmale):

Does this equation have a solution?

OpenStudy (anonymous):

yes ln(6)

OpenStudy (anonymous):

?

OpenStudy (mathmale):

Can you solve -6+e^x=0 for x? Yes. There is a solution. And therefore, the graph has a horiz. tan line. Now, please write out the original function.

OpenStudy (anonymous):

y=-6x+e^x

OpenStudy (mathmale):

Good. We found that there's a horiz. t. l. when x = ln 6. Please find f(ln 6)...evaluate the orignl fun. at x = ln 6.

OpenStudy (anonymous):

-6(ln6)+e^ln(6)

OpenStudy (mathmale):

Write the coordinates of this point in the form (x, y), please. Also recall that e^(ln 6) = 6.

OpenStudy (mathmale):

e^x and ln x are inverse functions.

OpenStudy (anonymous):

y=6-6*ln6 so ( ln (6) , 6-6ln(6) ) ( x , y )

OpenStudy (mathmale):

There is a horiz. tan line to the graph of f(x)=-6x+e^x at the point you have given. Does this match the answer?

OpenStudy (mathmale):

Remember: the correct answer, once we've determined that there is a horiz. tan. line, are the coordinates of the point of tangency.

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

*yes it does match my answer

OpenStudy (mathmale):

Cool. thanks for your perseverance. All done.

OpenStudy (anonymous):

Thank you for your help

OpenStudy (mathmale):

I wanted to emphasize that "the derivative at any point on a curve is the slope of the tangent line to the curve at that point." You're welcome!